# Equi-RO: A 4D mmWave Radar Odometry via Equivariant Networks

### v1

@@ Abstract
Autonomous vehicles and robots rely on accurate odometry estimation in GPS-denied environments. While LiDARs and cameras struggle under extreme weather, 4D mmWave radar emerges as a robust alternative with all-weather operability and velocity measurement. In this paper, we introduce Equi-RO, an equivariant network-based framework for 4D radar odometry. Our algorithm pre-processes Doppler velocity into invariant node and edge features in the graph, and employs separate networks for equivariant and invariant feature processing. A graph-based architecture enhances feature aggregation in sparse radar data, improving inter-frame correspondence. Experiments on the open-source dataset and self-collected dataset show Equi-RO outperforms state-of-the-art algorithms in accuracy and robustness. Overall, our method achieves 10.7% and 20.0% relative improvements in translation and rotation accuracy, respectively, compared to the best baseline on the open-source dataset.
@@ I Introduction
Odometry is indispensable for autonomous vehicles and robots, providing accurate support for downstream perception, planning, and control. When Global Positioning System (GPS) signals are degraded, such as under bridges, in tunnels, or during severe rain or snow, it becomes crucial to obtain relative localization using sensors such as LiDARs and cameras.
However, the performance of LiDARs and cameras degrades significantly under extreme weather conditions, limiting their applicability for odometry estimation in such scenarios. Recently, 4D millimeter-wave (mmWave) radar—characterized by compact size, cost efficiency, all-weather adaptability, velocity-measuring capability, and long detection range—has been increasingly regarded as a promising solution for robust perception and localization in autonomous driving, particularly under adverse weather conditions [1].
Current 4D mmWave radar odometry algorithms are mostly adapted from traditional LiDAR odometry methods. Some exploit unique properties of 4D mmWave radar point clouds like Doppler velocity and Radar Cross Section (RCS) to perform ego-velocity estimation and point cloud registration [2, 3]. Regarding learning-based approaches, a few pioneering works have adapted Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) for stand-alone 4D mmWave radar [4] and radar-camera fusion [5] odometry.
@@
Figure 1: Our Equi-RO method incorporates Doppler velocity compensation and graph construction to preserve noise-resilient and rotation-robust features. The derived equivariant and invariant features are fed into an equivariant Graph Neural Network (GNN)-based framework to obtain consistent and generalizable point cloud registration, yielding more accurate odometry results, especially in large rotations. Trajectory comparison on a partial segment of the loop3 split from the NTU4DRadLM dataset[6], with initialization aligned to the ground truth at the segment’s start, against [2].
@@
In learning-based odometry, accurate feature extraction and matching are critical steps. However, features in adjacent frames may become highly inconsistent when the ego-vehicle moves rapidly or turns sharply. Equivariant neural networks have emerged as a promising approach to address this issue [7, 8]. They learn features that transform predictably under specific geometric transformations, such as 3D translations and rotations represented by rigid-body motion group SE(3), thereby enhancing feature matching performance. The geometric symmetry can improve the generalizability and interpretability of learning-based odometry systems.
@@
The 4D mmWave radar provides additional physical attributes compared to LiDAR point clouds. For example, it measures the RCS of each point, which can be regarded as SE(3)-invariant since it reflects the strength of the returned signal and is determined mainly by the target’s size, shape, and material rather than the sensor’s pose. In addition, the Doppler effect enables estimation of the radial relative velocity of detected targets. However, Doppler velocity is neither equivariant nor invariant under SE(3) transformations, as it depends on both the sensor’s motion and the target’s motion, which limits the applicability of existing equivariant networks to 4D mmWave radar odometry. Moreover, 4D mmWave radar point clouds are inherently noisier, sparser, and more irregular than those captured by scanning LiDARs. Consequently, effectively aggregating point-wise features with those of their spatial neighbors remains a significant challenge for learning-based 4D mmWave radar odometry.
@@
In this paper, we propose Equi-RO, a novel framework for 4D mmWave radar odometry built upon equivariant neural networks. Compared with existing 4D mmWave radar odometry algorithms, we design a pre-processing pipeline to derive invariant Doppler features and construct a graph representation to mitigate noise and sparsity. The equivariant and invariant features are then processed through separate networks to estimate the relative transformation. The invariant velocity features and equivariant network design enable our method to remain robust in challenging real-world driving scenarios, where existing methods often degrade. Leveraging both a public dataset [6] and a self-collected dataset, we demonstrate consistent improvements over state-of-the-art 4D mmWave radar odometry approaches.
@@
    We develop a novel 4D mmWave radar odometry algorithm based on equivariant neural networks, capable of producing accurate and robust odometry estimates under large rotations and fast motions, as commonly encountered in sharp turns and challenging driving maneuvers.
@@
    We design a dedicated pre-processing pipeline for Doppler velocity that derives invariant features, thereby stabilizing motion-related features. And we design a framework that separately extracts equivariant and invariant features to preserve geometric consistency, making it applicable to 4D mmWave radar point clouds with physical attributes.
@@
    We conduct extensive experiments on both an open-source dataset and a self-collected dataset, along with ablation studies, to validate the effectiveness of the proposed method across diverse scenarios.
  4. 4.
    The source code will be released after receiving the final decision.
@@ II Related Works
In this section, we briefly review existing algorithms for 4D mmWave radar odometry, followed by a discussion of equivariant neural networks.
II-A 4D mmWave Radar Odometry
Existing 4D mmWave radar odometry approaches can be broadly categorized into traditional and learning-based methods.
Traditional odometry methods, which do not involve neural networks, primarily address two challenges: (1) the effective utilization of Doppler velocity and RCS information, and (2) the handling of noise and sparsity in radar point clouds. For the first aspect, Doppler velocity is commonly employed for ego-velocity estimation [9] and dynamic object removal [3], while RCS is often used for point cloud filtering [10] and weighted point matching [11]. In the latter case, some works have proposed to use scan-to-submap registration [3] and adaptive probability-based registration [2, 12] to handle the noise and sparsity.
Learning-based approaches replace traditional point cloud registration and pose regression with deep neural networks. As an early pioneer, Lu et al. [4] employed CNNs to extract features from radar point clouds. RaFlow [13] formulates odometry estimation by scene flow prediction. 4DRO-Net [14] and 4DRVO-Net [5] integrate velocity-aware attention cost volumes into a coarse-to-fine optimization framework to iteratively estimate poses. CAO-RONet [15] incorporates local completion and context-aware association to match noisy points and suppress outliers. To the best of our knowledge, although GNNs can explicitly aggregate features from spatially neighboring points, they have not yet been applied to 4D mmWave radar odometry.
II-B Equivariant Neural Networks
@@
Equivariant neural networks are designed to produce predictable changes in their outputs when specific transformations are applied to the inputs. This property enhances both the generalizability and interpretability of the models. Since Cohen and Welling first introduced the concept of equivariance into neural networks [7], various equivariant designs have been proposed for GNNs [16, 17]. For example, Satorras et al. [17] propose Equivariant Graph Neural Network (EGNN), which maintains equivariance to the Euclidean group E(n)—including translations, rotations, and reflections—by incorporating the relative squared distances between coordinates in the edge feature aggregation process.
@@
Significant efforts have also been devoted to designing equivariant architectures for point cloud analysis. Tensor Field Network (TFN) [18] employs spherical harmonics to construct learnable weight kernels, preserving equivariance to SE(3) for point cloud data. Vector Neurons (VN) [19] and Lie Neurons (LN) [20] extend traditional Multi-Layer Perceptrons (MLPs) to operate on vector features, achieving equivariance to 3D rotation group SO(3) and semisimple Lie groups, respectively. EPN [21] and E2PN [8] further introduce SE(3)-equivariant convolutional architectures for 3D point clouds. Despite these advances, the application of equivariant neural networks to 4D mmWave radar point clouds remains largely unexplored.
@@ III Methodology
This section provides a detailed description of our proposed Equi-RO algorithm. We first present an overview of the approach in Sec. III-A, followed by the pre-processing pipeline for 4D mmWave radar point clouds in Sec. III-B. Sec. III-C introduces the graph construction process and the definition of node and edge features, while Sec. III-D outlines the overall network architecture of Equi-RO.
@@ III-A Overview
The overall pipeline of Equi-RO is illustrated in Fig. 2. Two consecutive frames of 4D mmWave radar point clouds are first pre-processed to estimate the ego-vehicle’s velocity and to generate invariant compensated node and edge velocities based on Doppler velocity. The processed point clouds are represented as graphs, with carefully defined node and edge features. An SE(3)-equivariant network based on LN [20] is employed to extract and aggregate node-level equivariant features, while both node- and edge-level invariant features are fed into an EGNN-based [17] module. The combined features are ultimately used to estimate the relative transformation $(R,t)$ between the two input frames, achieving robust registration even under large rotations.
Figure 2: Overview of the Equi-RO algorithm. Doppler velocity is used to derive invariant node and edge features, which are combined with node and edge features within a unified graph-based network framework to robustly estimate the relative transformation $(R,t)$ between two consecutive radar frames.
@@ III-B Point Cloud Pre-processing
As mentioned above, 4D mmWave radar can directly measure the radial relative velocity of detected targets via the Doppler effect. In this section, we fully exploit the Doppler velocity to estimate the ego-vehicle’s velocity and to derive compensated node and edge velocities for subsequent use in the GNN-based equivariant framework.
@@ III-B1 Ego-velocity Estimation
We apply an iterative reweighted least squares algorithm [22] to estimate the ego-velocity $\mathbf{v}_{ego}$. The Doppler velocity $v_{i}^{dop}$ of each point $i$ is affected by the coordinate $\bm{\rho}_{i}$ of this point in the sensor frame. In this work, Doppler velocity is defined as positive for approaching targets and negative for receding ones. For a point with absolute velocity $\mathbf{v}_{i}^{abs}$, its Doppler velocity $v_{i}^{dop}$ can be represented as:
@@
where $\bm{\rho}_{i}^{T}$ is the transpose of $\bm{\rho}_{i}$. Assuming that most points in the scene are stationary, and assigning each point $i$ a weight $\omega_{i}$ indicating its likelihood of being stationary, we formulate the following weighted least squares problem for a point cloud with $N$ points:
$\underset{\mathbf{v}_{ego}}{\min}\sum_{i=1}^{N}\omega_{i}\left\|v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}\right\|.$ (2)
Here, the weights are iteratively updated as $\omega_{i}=1/(v_{r}+\epsilon)$, where $v_{r}=\left\|v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}\right\|$ and $\epsilon=1\times 10^{-5}$. Upon convergence, $\mathbf{v}_{ego}$ provides a robust estimate of the ego-vehicle’s velocity.
@@ III-B2 Node Velocity Compensation
Once $\mathbf{v}_{ego}$ is estimated, the compensated node velocity $v_{i}^{dop^{\prime}}$ for each point $i$ can be computed as:
$v_{i}^{dop^{{}^{\prime}}}=v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}.$ (3)
As illustrated in Fig. 3, $v_{i}^{dop^{\prime}}$ corresponds to $\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{i}^{abs}$, i.e., the radial component of the absolute velocity $\mathbf{v}_{i}^{abs}$ of point $i$. Notably, $v_{i}^{dop^{\prime}}$ is invariant under SO(3) transformations of the radar, providing a stable feature representation for point cloud registration under large rotational variations.
Since $\mathbf{v}_{ego}$ has already been obtained to enhance relative translation estimation, incorporating the SO(3)-invariant $v_{i}^{dop^{\prime}}$ further strengthens the robustness of feature matching, thereby improving rotational alignment accuracy.
@@ III-B3 Edge Velocity Compensation
As Fig. 3 presents, when $\mathbf{v}_{i}^{abs}=\mathbf{v}_{j}^{abs}$, $v_{ij}$ can be further expressed as:
$v_{ij}=\mathbf{v}_{i}^{abs}\cdot\frac{(\bm{\rho}_{i}-\bm{\rho}_{j})}{\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|}=\mathbf{v}_{i}^{abs}\cdot\frac{\bm{\rho}_{r}}{\left\|\bm{\rho}_{r}\right\|},$ (5)
where $\bm{\rho}_{r}$ denotes the relative coordinates between $i$ and $j$.
It is evident that $v_{ij}$ remains invariant under any rotation or translation of the sensor. Therefore, if both nodes of an edge share the same absolute velocity—which holds when they belong to the same rigid object with no angular velocity—the compensated edge relative velocity $v_{ij}$ is SE(3)-invariant. Moreover, even when the two nodes have different absolute velocities, $v_{ij}$ is still SO(3)-invariant, since both $v_{i}^{dop^{\prime}}$ and $v_{j}^{dop^{\prime}}$ are individually SO(3)-invariant.
Figure 3: Illustrations of (a) node and (b) edge velocity compensation.

### v2

@@ Abstract
Autonomous vehicles and robots rely on accurate odometry estimation in GPS-denied environments. While LiDARs and cameras struggle under extreme weather, 4D mmWave radar emerges as a robust alternative with all-weather operability and velocity measurement. In this paper, we introduce Equi-RO, an equivariant network-based framework for 4D radar odometry. Our algorithm pre-processes Doppler velocity into invariant node and edge features in the graph, and employs separate networks for equivariant and invariant feature processing. A graph-based architecture enhances feature aggregation in sparse radar data, improving inter-frame correspondence. Experiments on an open-source dataset and a self-collected dataset show Equi-RO outperforms state-of-the-art algorithms in accuracy and robustness. Overall, our method achieves 10.7% and 13.4% relative improvements in translation and rotation accuracy, respectively, compared to the best baseline on the open-source dataset.
@@ I Introduction
Odometry is indispensable for autonomous vehicles and robots, supporting downstream perception, planning, and control. When Global Positioning System (GPS) signals are degraded, localization relies primarily on odometry from onboard sensors such as LiDARs and cameras. However, these sensors often struggle in extreme weather, whereas 4D millimeter-wave (mmWave) radar (4D radar) provides compact, cost-efficient, all-weather, velocity-measuring, and long-range sensing, making it a promising alternative for robust perception and localization under adverse conditions [12]. Current 4D radar odometry mostly adapts traditional LiDAR methods and exploits Doppler velocity and Radar Cross Section (RCS) for ego-velocity estimation and point cloud registration [35, 38], while a few learning-based works employ convolutional and recurrent neural networks for 4D radar [21] and radar-camera fusion [39] odometry.
@@
Figure 1: Our Equi-RO method incorporates Doppler velocity compensation and graph construction to preserve noise-resilient and rotation-robust features. The derived equivariant and invariant features are fed into an equivariant Graph Neural Network (GNN)-based framework for consistent and generalizable point cloud registration, yielding more accurate odometry results, especially in large rotations. Trajectory comparison on a partial segment of the loop3 split from the NTU4DRadLM dataset [34], where both methods are initialized to the ground truth at the segment’s start, comparing against [35].
@@
In learning-based odometry, feature extraction and matching are critical. However, features in adjacent frames may become highly inconsistent when the ego-vehicle moves rapidly or turns sharply. Equivariant neural networks have emerged as a promising approach to address this issue [5, 37]. They learn features that transform predictably under, e.g., rigid body transformation group SE(3) (equivariance property), or remain fixed under such transformations (invariance property).
@@
In 4D radar sensing, the Doppler effect enables estimation of targets’ radial relative velocity. However, Doppler velocity is neither equivariant nor invariant under the sensor’s transformations, as it depends on both the sensor’s and target’s motion, limiting the applicability of equivariant networks to 4D radar odometry. Moreover, 4D radar point clouds, with typically only thousands of points per frame, are noisier, sparser, and more irregular than LiDAR point clouds. Consequently, effectively aggregating point-wise features from spatial neighbors remains a challenge for learning-based 4D radar odometry.
@@
In this paper, we propose Equi-RO, a novel 4D radar odometry framework built upon equivariant networks (Fig. 1). We design a pre-processing pipeline to derive invariant Doppler features and construct a graph from the point cloud to mitigate noise and sparsity. Compared with raw point sets or voxel grids, a graph representation explicitly connects spatially and kinematically consistent neighbors, which can suppress isolated noise and stabilize feature extraction for 4D radar point clouds. The invariant velocity features and equivariant network design enable our method to remain robust in challenging real-world driving scenarios, where existing methods often degrade. Leveraging both a public dataset [34] and a self-collected dataset, we demonstrate consistent improvements over state-of-the-art 4D radar odometry approaches.
@@
    We develop a novel 4D radar odometry algorithm based on equivariant neural networks, providing accurate and robust odometry estimates under large rotations, as commonly encountered in sharp turns and challenging driving maneuvers.
@@
    We design a dedicated pre-processing pipeline for Doppler velocity that derives invariant features, thereby stabilizing motion-related features. We also design a framework that separately extracts equivariant and invariant features to preserve geometric consistency, making it applicable to 4D radar point clouds with physical attributes.
@@
    We conduct extensive experiments on both an open-source dataset and a self-collected dataset, along with ablation studies, to validate the effectiveness of the proposed method across diverse scenarios. The project is available at: https://github.com/hanzy21/Equi-RO.
@@ II Related Works
In this section, we briefly review existing algorithms for 3D point cloud odometry and 4D radar odometry, followed by a discussion of equivariant neural networks.
II-A 3D Point Cloud Odometry
3D point cloud odometry for LiDAR and RGB-D cameras can be categorized into traditional and learning-based methods. Traditional methods mainly utilize geometric registration, such as Iterative Closest Point (ICP) [1] and Normal Distributions Transform (NDT) [2]. Subsequent Generalized-ICP (GICP) [27] and KISS-ICP [31] further improve robustness through local surface modeling. Learning-based methods such as FCGF [4] and DeepVCP [23] replace handcrafted features with learned correspondence models on dense point clouds.
Despite their success on dense point clouds, these methods’ heavy reliance on dense and smooth geometry limits their applicability to sparse and noisy data from 4D radars.
II-B 4D Radar Odometry
Several core ideas of 3D point cloud odometry including geometric registration and local feature extraction are transferable to 4D radar odometry. However, leveraging radar-specific measurements (Doppler and RCS) and handling the sparsity and noise of radar point clouds remain open challenges. These issues have led to recent works on 4D radar odometry.
For measurements utilization, Doppler velocity is commonly employed for ego-velocity estimation [8, 32], dynamic object removal [38, 14] and velocity-aware attention module design [22, 39], while RCS is often used for point cloud filtering [15, 13] and weighted point matching [16] since it indicates the intensity of the point.
As for handling the sparsity and noise of radar point clouds, classical 4D radar odometry uses scan-to-submap [38, 17], adaptive probability-based [35, 18], and voxel-based [28] registration to improve the robustness. For learning-based methods, RaFlow [7] predicts scene flow, while CAO-RONet [19] incorporates local completion to match noisy points.
To the best of our knowledge, although GNNs can explicitly aggregate features from spatially neighboring points, they have not yet been applied to 4D radar odometry.
II-C Equivariant Neural Networks
@@
Equivariant neural networks produce predictable outputs under specific input transformations, enhancing generalizability and interpretability. Since Cohen and Welling introduced group equivariance into neural networks [5], various equivariant designs have been proposed for GNNs [9, 26]. For example, Equivariant Graph Neural Network (EGNN) [26] maintains equivariance to the Euclidean group E(n), which includes translations, rotations, and reflections, by incorporating the relative distances in the edge feature aggregation.
@@
For point cloud analysis, Tensor Field Network [30] achieves SE(3)-equivariance via spherical harmonics, while Vector Neurons (VN) [6] and Lie Neurons (LN) [20] extend traditional Multi-Layer Perceptrons (MLPs) to vector features, achieving equivariance to 3D rotation group SO(3) and semi-simple Lie groups, respectively. EPN [3] and E2PN [37] further introduce SE(3)-equivariant convolutional architectures for 3D point clouds. However, equivariant models remain largely underexplored for 4D radar point clouds.
@@ III Methodology
This section provides a detailed description of our proposed Equi-RO algorithm.
@@ III-A Overview
Fig. 2 shows the pipeline of Equi-RO. Standard 4D radar odometry faces two primary challenges: the sparsity of radar point clouds and the ego-motion dependent nature of Doppler velocities. To address sparsity and noise, we structure the point cloud as a graph rather than isolated points. This allows us to explicitly define edge features that capture local relative patterns, aggregating information from neighbors to improve robustness. To address the motion coupling of Doppler velocity, we mathematically derive invariant compensated velocities to decouple the object’s motion from the ego-vehicle’s rotation.
Based on these pre-processed inputs, we extract equivariant geometric structures (e.g., 3D coordinates that transform consistently with the sensor) and invariant physical features (e.g., RCS and our derived Doppler velocity that remain unchanged under the sensor’s rotation or translation), and process these features with different networks. The equivariant features enable the network to reason about relative motion during scan matching, while the invariant features remain consistent across different sensor orientations, providing pose-independent cues. Specifically, an SO(3)-equivariant network based on LN [20] is employed to extract node-level equivariant features, while both node- and edge-level invariant features are fed into an EGNN-based [26] module. The output features are used to estimate the relative transformation $(\bm{R},\bm{t})$ between the two input frames, achieving robust registration even under large rotations.
Figure 2: Overview of the Equi-RO algorithm. Doppler velocity is used to derive invariant node and edge features, which are combined with node and edge features within a unified graph-based network framework to robustly estimate the relative transformation $(\bm{R},\bm{t})$ between two consecutive radar frames.
@@ III-B Point Cloud Pre-processing
As mentioned above, 4D radar can directly measure the radial relative velocity of targets via the Doppler effect. In this section, we fully exploit the Doppler velocity to estimate the ego-velocity and to derive compensated node and edge velocities for subsequent GNN-based equivariant framework.
@@ III-B1 Ego-velocity Estimation
We apply an iterative reweighted least squares algorithm [33] to estimate the ego-velocity $\mathbf{v}_{ego}$. For a point $i$ with absolute velocity $\mathbf{v}_{i}^{abs}$, its Doppler velocity $v_{i}^{dop}$ can be represented as:
@@
where $\bm{\rho}_{i}$ is the coordinate of this point. We define the Doppler residual $v_{i}^{r}=\left\|v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}\right\|$, which should be zero for stationary points. To robustly suppress moving points, we can formulate the following weighted least squares problem for a point cloud with $N$ points:
$\underset{\mathbf{v}_{ego}}{\min}\sum_{i=1}^{N}\omega_{i}{v_{i}^{r}}^{2}.$ (2)
The weight $\omega_{i}$ indicates the likelihood of a point being stationary and is iteratively updated as $\omega_{i}=1/(v_{i}^{r}+\epsilon)$, with $\epsilon=1\times 10^{-5}$. Upon convergence, $\mathbf{v}_{ego}$ yields a robust ego-velocity estimate.
@@ III-B2 Node Velocity Compensation
Once $\mathbf{v}_{ego}$ is estimated, the compensated node velocity $v_{i}^{dop^{\prime}}$ for each point $i$ is:
$v_{i}^{dop^{{}^{\prime}}}=v_{i}^{dop}-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{ego}=-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{i}^{abs}.$ (3)
As illustrated in Fig. 3 (left), $v_{i}^{dop^{\prime}}$ is the radial component of the absolute velocity $\mathbf{v}_{i}^{abs}$ of point $i$. If the radar coordinate frame is rotated by a matrix $\bm{R}\in SO(3)$ with $\bm{R}^{T}\bm{R}=\bm{I}$, $v_{i}^{dop^{\prime}}$ remains the same as:
$v_{i}^{dop^{{}^{\prime\prime}}}$ $=-\frac{(\bm{R\rho}_{i})^{T}}{\left\|\bm{R\rho}_{i}\right\|}\cdot\bm{R}\mathbf{v}_{i}^{abs}=-\frac{\bm{\rho}_{i}^{T}\bm{R}^{T}\bm{R}\mathbf{v}_{i}^{abs}}{\left\|\bm{\rho}_{i}\right\|}$
$=-\frac{\bm{\rho}_{i}^{T}}{\left\|\bm{\rho}_{i}\right\|}\cdot\mathbf{v}_{i}^{abs}=v_{i}^{dop^{{}^{\prime}}}.\vskip-14.22636pt$ (4)
Therefore, $v_{i}^{dop^{\prime}}$ is invariant under SO(3) transformations of the radar, providing a stable feature representation for point cloud registration under large rotational variations.
@@ III-B3 Edge Velocity Compensation
Since both $v_{i}^{dop^{\prime}}$ and $v_{j}^{dop^{\prime}}$ are individually SO(3)-invariant, and rotations preserve vector norms and pairwise distances, it is evident that $v_{ij}$ is also SO(3)-invariant.
Moreover, as Fig. 3 (right) presents, if both nodes of an edge have identical absolute velocities, i.e., $\mathbf{v}_{i}^{abs}=\mathbf{v}_{j}^{abs}$, we can substitute Eq. (3) into Eq. (5), $v_{ij}$ as:
$v_{ij}=\frac{(\bm{\rho}_{j}^{T}-\bm{\rho}_{i}^{T})}{\left\|\bm{\rho}_{i}-\bm{\rho}_{j}\right\|}\cdot\mathbf{v}_{i}^{abs}=\frac{\bm{\rho}_{r}^{T}}{\left\|\bm{\rho}_{r}\right\|}\cdot\mathbf{v}_{i}^{abs},$ (6)
where $\bm{\rho}_{r}$ denotes the relative coordinates from $i$ to $j$. If the radar undergoes a rigid-body transformation $(\bm{R},\bm{t})\in SE(3)$, the translation $\bm{t}$ does not change $\mathbf{v}_{i}^{abs}$ and $\bm{\rho_{r}}$. So we have:
$v_{ij}^{{}^{\prime}}$ $=\frac{(\bm{R\rho}_{r})^{T}}{\left\|\bm{R\rho}_{r}\right\|}\cdot(\bm{R}\mathbf{v}_{i}^{abs})=\frac{\bm{\rho}_{r}^{T}\bm{R}^{T}\bm{R}\mathbf{v}_{i}^{abs}}{\left\|\bm{\rho}_{r}\right\|}$ (7)
$=\frac{\bm{\rho}_{r}^{T}}{\left\|\bm{\rho}_{r}\right\|}\cdot\mathbf{v}_{i}^{abs}=v_{ij}.$
Hence, the edge compensated velocity $v_{ij}$ is SE(3)-invariant when its two nodes share the same velocity, which holds when they belong to a rigid object with no angular velocity.
Figure 3: Illustrations of (left) node and (right) edge velocity compensation. The yellow arrow in the left figure indicates the compensated node velocity, which remains SO(3)-invariant. The yellow arrow in the right figure is the compensated edge velocity when the two nodes share the same absolute velocity, which is SE(3)-invariant.